by selling a ball for rs 39, a shopkeeper gains 30%.at what price should he sell it to gain 40%.

Grade:8

1 Answers

Anshika Pattnaik
6 years ago
Let cost price be xProfit = cost price * profit% = x * 30/100Selling price = cost price + profit = x + 30x/100 = 13x/10Acc. To ques.Selling price = 39 rs. 13x /10 =39x = 30rs.For 40% profit in selling price Cost price + 40%cost price 30 + 40*30/10030+1242rs.